# A1: Asynchronous Test-Time Scaling via Conformal Prediction

### v1

@@ 3.3 How to perform rejection sampling via conformal prediction?
Given input tokens $x_{1:n-1}$, the draft model proposes $m$ candidate continuations of length $K_{d}$, denoted $\tilde{y}_{n:n+K_{d}-1}^{j}$, by sampling from the same draft distribution $q_{d}$:
@@
with the joint specified tokenwise as
$q_{d}\!\left(\tilde{y}_{n:n+K_{d}-1}^{j}\mid x_{1:n-1}\right)=\prod_{i=0}^{K_{d}-1}q_{d}\!\left(\tilde{y}_{n+i}^{j}\,\middle|\,x_{1:n-1},\tilde{y}_{n:n+i-1}^{j}\right),\quad j=1,\dots,m.$ (16)
@@
For each candidate sampling $\tilde{y}_{n:n+K_{d}-1}^{j}$, we score it under the target model $q_{t}$ by computing the logits/likelihood $q_{t}(\tilde{y}_{n:n+K_{d}-1}^{j}\mid x_{1:n-1})$ and converting to a conformity score. Using the calibration set, we compute a $p$-value for each candidate and reject those inside $\mathcal{C}_{\alpha}$; otherwise accept.
@@
Although classical rejection sampling discards the rejected samples from the draft model and resamples entirely from the target model, to save the token budget we proceed as follows. In each turn, the per-round target-side token budget is $K_{t}$: for each accepted candidate in $\mathcal{C}_{\alpha}$, we let the target model $q_{t}$ continue generation using that candidate as a prefix for up to $K_{t}$ tokens, stopping earlier if an end token is encountered.
@@
We iterate the above stages until a final answer is detected, the maximum number of turns is reached, or the overall token limit is exceeded. As highlighted in Fig. 5, increasing the number of turns enables sequential test-time scaling (blue box), while increasing the number of candidates per round enables parallel test-time scaling (green box).
Table 1: The comparison when the draft and target models are from different families, under the setting on marginal coverage (Mar Cov.) and conditional coverage (Con Cov.). indicates that both are reasoning models, whereas indicates that only the target model is a reasoning model. The results highlighted in red indicate the lossless acceleration method.
Dataset Draft Model (DM) / Target Model (TM) Type Mar Cov. Con Cov. DM Baseline TM Baseline
MATH100
DeepSeek-R1-Distill-Qwen-1.5B / QwQ-32B 87.0 94.0 83.0 96.0
Qwen2.5-7B-Instruction / QwQ-32B 86.0 96.0 84.0 96.0
Llama-3.1-8B-Instruct / QwQ-32B 87.0 95.0 75.0 96.0
DeepSeek-R1-Distill-Llama-8B / QwQ-32B 96.0 96.0 85.0 96.0
DeepSeek-R1-Distill-Qwen-1.5B / s1.1-32B 86.0 95.0 83.0 96.0
Qwen2.5-7B-Instruction / s1.1-32B 89.0 96.0 84.0 96.0
Llama-3.1-8B-Instruct / s1.1-32B 79.0 85.0 75.0 96.0
DeepSeek Llama-3.1-8B-Instruct / s1.1-32B 88.0 95.0 85.0 96.0
AIME24
DeepSeek-R1-Distill-Qwen-1.5B / QwQ-32B 86.7 66.7 60.0 86.7
Qwen2.5-7B-Instruction / QwQ-32B 46.7 33.3 33.3 86.7
Llama-3.1-8B-Instruct / QwQ-32B 33.3 40.0 13.3 86.7
DeepSeek Llama-3.1-8B-Instruct / QwQ-32B 80.0 80.0 80.0 86.7
DeepSeek-R1-Distill-Qwen-1.5B / s1.1-32B 73.3 80.0 60.0 86.7
Qwen2.5-7B-Instruction / s1.1-32B 40.0 33.3 33.3 86.7
Llama-3.1-8B-Instruct / s1.1-32B 26.7 40.0 13.3 86.7
DeepSeek Llama-3.1-8B-Instruct / s1.1-32B 73.3 73.3 80.0 86.7
AIME25
DeepSeek-R1-Distill-Qwen-1.5B / QwQ-32B 53.3 46.7 40.0 73.3
Qwen2.5-7B-Instruction / QwQ-32B 33.3 40.0 26.7 73.3
Llama-3.1-8B-Instruct / QwQ-32B 26.7 33.3 20.0 73.3
DeepSeek Llama-3.1-8B-Instruct / QwQ-32B 66.7 60.0 46.7 73.3
DeepSeek-R1-Distill-Qwen-1.5B / s1.1-32B 60.0 60.0 40.0 66.7
Qwen2.5-7B-Instruction / s1.1-32B 33.3 40.0 26.7 66.7
Llama-3.1-8B-Instruct / s1.1-32B 26.7 20.0 20.0 66.7
DeepSeek Llama-3.1-8B-Instruct / s1.1-32B 66.7 60.0 46.7 66.7
AMC23
DeepSeek-R1-Distill-Qwen-1.5B / QwQ-32B 88.0 90.0 74.0 94.0
Qwen2.5-7B-Instruction / QwQ-32B 76.0 72.0 68.0 94.0
Llama-3.1-8B-Instruct / QwQ-32B 68.0 68.0 44.0 94.0
DeepSeek Llama-3.1-8B-Instruct / QwQ-32B 92.0 94.0 80.0 94.0
DeepSeek-R1-Distill-Qwen-1.5B / s1.1-32B 86.0 86.0 74.0 96.0
Qwen2.5-7B-Instruction / s1.1-32B 74.0 78.0 68.0 96.0
Llama-3.1-8B-Instruct / s1.1-32B 54.0 64.0 44.0 96.0
DeepSeek Llama-3.1-8B-Instruct / s1.1-32B 92.0 93.0 80.0 96.0
@@ 4 Experiment
We provide the hyperparameters used in our experiments in Appendix 4. Table 1 reports the results of asynchronous test-time scaling when the draft model (DM) and target model (TM) come from different families. We evaluate performance under two distinct settings: marginal coverage (Mar Acc.), which measures whether the predefined budget and prediction set align on average across test inputs, and conditional coverage (Con Acc.), which imposes a stricter requirement that the guarantee holds for each individual input instance. We have the following key takeaways: i) Speculative decoding can match the performance of the target model itself. This approach effectively reduces computational overhead while maintaining high-quality outputs, offering a feasible solution for scaling inference. ii) The most challenging datasets, AIME24/25, show strong performance in marginal coverage setup, while the other two datasets (MATH100 and AMC23) demonstrate superior results in conditional coverage setup. This highlights that while marginal coverage allocates more computational resources to the most difficult parts of the tasks effectively, conditional coverage ensures more reliable results at the individual input level, especially in simpler tasks, ensuring that each question is answered correctly. iii) This shows that even though reasoning models typically consume more tokens during inference, when constrained to the same token budget, using a reasoning model as the draft model yields stronger scaling performance compared to a non-reasoning model.
Table 2: The comparison when the draft and target models are from the same families.
@@
SpecThink
MATH100 Qwen2.5-7B/32B-Instruct 86.0 81.0 73.7 79.0
s1.1-7B/32B 88.0 87.0 73.7 85.0
DeepSeek-R1-Distill-Qwen-1.5B/32B 88.0 87.0 76.7 84.0
Skywork-OR1-7B/32B 88.0 89.0 75.8 70.0
AIME24 Qwen2.5-7B/32B-Instruct 33.3 40.0 33.3 33.3
s1.1-7B/32B 66.7 73.3 26.7 40.0
DeepSeek-R1-Distill-Qwen-1.5B/32B 86.7 80.0 66.7 66.7
Skywork-OR1-7B/32B 86.7 80.0 73.3 60.0
AIME25 Qwen2.5-7B/32B-Instruct 40.0 33.3 40.0 26.7
s1.1-7B/32B 53.3 53.3 40.0 33.3
DeepSeek-R1-Distill-Qwen-1.5B/32B 60.0 53.3 35.7 46.7
Skywork-OR1-7B/32B 60.0 53.3 53.3 40.0
AMC23 Qwen2.5-7B/32B-Instruct 72.0 70.0 72.0 74.0
s1.1-7B/32B 82.0 78.0 78.0 76.0
DeepSeek-R1-Distill-Qwen-1.5B/32B 92.0 88.0 80.0 82.0
Skywork-OR1-7B/32B 96.0 94.0 86.0 82.0
@@ 4.2 Performance of Budget prediction
In this section, we present the accuracy of budget prediction under the settings of marginal coverage and conditional coverage. It demonstrates whether our method can accurately control the number of target model interventions in rejection sampling. This directly shows the accuracy of conformal prediction in estimating the rejection rate.
@@
In Figure 6(b), we report the accuracy of the target-model intervention rate under marginal coverage, where the rejection rate is predicted at the dataset level. Budget-prediction accuracy across the full dataset is high—especially with the 64-sample configuration, whose absolute error stays within $5\%$. With $K_{d}=500$ tokens for calibration and $K_{d}=500$ for sampling, the error remains within $2\%$. This directly highlights the importance of constructing a diverse calibration set for maintaining high conformal-prediction accuracy.
@@
In practice, we require precise per-batch budget control—rather than a single aggregate budget over the entire test set. Because the target-model server has limited capacity and cannot process all requests concurrently, inference is executed in batches, and the token budget must be enforced for each batch to satisfy the load constraint. In Figure 6(a), we report the accuracy of the target model intervention rate under conditional coverage. Under online calibration with a rejection rate of $25\%$, when $K_{d}=500$ in both the calibration and sampling stages, the 16-sample and 64-sample settings achieve similar accuracy. By contrast, when the calibration stage uses $K_{d}=500$ but the sampling stage uses $K_{d}=700$, the 64-sample setting attains significantly higher budget-prediction accuracy. This indicates that increasing the number of parallel samples can improve budget-prediction accuracy when the sampling token budget differs from the calibration token budget (i.e., under a calibration–sampling token-budget mismatch).
@@ 4.3 Asynchronous Test-time Scaling within The Same Model Families
In this section, we examine the performance across models within the same family, including both reasoning and non-reasoning variants in Table 2. In this setting, since the target model and draft models share the same vocabulary, we can compare against baselines that are only applicable to models within the same family, such as Speculative Thinking (Yang et al., 2025), denoted as SpecThink. Our findings are as follows: i): When the draft and target models belong to the same family, larger target models consistently outperform their smaller counterparts, confirming the benefit of intra-family consistency. ii): Across datasets, DeepSeek and Skywork show the strongest gains on challenging benchmarks (AIME, AMC), while Qwen2.5 performs competitively on MATH100 but lags significantly on harder tasks. iii): Moreover, s1.1 achieves moderate improvements, usually surpassing Qwen but not reaching the level of Skywork or DeepSeek. iv): Finally, speculative baselines (SpecReason and SpecThink) generally underperform compared with target model, suggesting that their effectiveness remains limited in reasoning-intensive settings.
Figure 7: Token budget at a maximum length of 20000 and 16 samples.
@@ 4.4 Analysis of token budget and latency
In this part, we analyze the trade-offs between latency and throughput under different sampling and token budget settings. i): As shown in Figure 8(a), the latency of SpecReason and SpecThinking consistently increases with the number of samples, highlighting the cost of scaling up sampling. In contrast, our method significantly reduces the sampling latency and achieves the lowest inference latency under the condition coverage setting. ii): The overhead of online calibration is nearly negligible, particularly at larger sample sizes. iii): Meanwhile, Figure 8(b) illustrates how throughput varies with the per-sample token budget for methods such as SpecReason and SpecThinking, revealing diminishing returns as the token budget becomes large. Our method is able to maintain high throughput even under very large budgets, especially in the marginal coverage setting. Overall, these results (Figure 8) provide insights into the balance between efficiency and performance when designing inference strategies. Under the setting of 16 samples and a maximum token budget of 8192 per sample, we achieved a 56.7x speedup in inference and a 4.14x throughput improvement compared to the baseline.
@@
We present in Figure 7 the token consumption under the 16-sample setting. This includes sampling performed solely by the target model or the draft model; as well as asynchronous sampling under both condition coverage and marginal coverage settings. Compared with the two baselines of the draft model and the target model, our method can significantly reduce token consumption, especially under the condition coverage setting, since it enables budget prediction at the instance level.
@@
We presented A1 (Asynchronous test-time scaling with conformal prediction), a framework that addresses the core inefficiencies of test-time scaling in LLMs. By refining arithmetic intensity and introducing online calibration with a rejection sampling pipeline, A1 effectively controls rejection rates while reducing latency and memory overhead. Experiments on multiple reasoning benchmarks confirm that A1 achieves better efficiency and reliability than speculative baselines. This work establishes A1 as a practical and principled approach for scalable LLM inference, with potential extensions to dynamic adaptation and real-world deployment.

### v2

@@ 3.3 How to perform rejection sampling via conformal prediction?
Given input tokens $x_{1:n-1}$, the draft model proposes $m$ candidate continuations of length $K_{d}$ in each turn, denoted $\tilde{y}_{n:n+K_{d}-1}^{j}$, by sampling from the draft model $q_{d}$:
@@
For each candidate sampling $\tilde{y}_{n:n+K_{d}-1}^{j}$, we score it under the target model $q_{t}$ by computing the logits $q_{t}(\tilde{y}_{n:n+K_{d}-1}^{j}\mid x_{1:n-1})$ and converting to a conformity score. Using the calibration set, we compute a $p$-value for each candidate and reject those inside $\mathcal{C}_{\alpha}$; otherwise accept.
@@
Although classical rejection sampling discards the rejected samples from the draft model and resamples entirely from the target model, to save the token budget we proceed as follows. In each turn, the per-turn target-side token budget is $K_{t}$: for each candidate in $\mathcal{C}_{\alpha}$, we let the target model $q_{t}$ continue generation using that candidate from $q_{d}$ as a prefix for up to $K_{t}$ tokens, stopping earlier if an end token is encountered. In Appendix A.4, we provide a comparison between continuing sampling and resampling with $q_{t}$ for large-scale scaling performance.
@@
We iterate the above rejection sampling until a final answer is detected, the maximum number of turns is reached, or the overall token limit is exceeded. As highlighted in Fig. 5, increasing the number of turns enables sequential test-time scaling (blue box), while increasing the number of candidates per turn enables parallel test-time scaling (green box).
Table 1: Comparison when draft and target models are from different families under marginal (Mar Cov.) and conditional (Con Cov.) coverage. $\bm{S_{\text{Mar}}}$, $\bm{S_{\text{Con}}}$ = end-to-end speedup (×) under marginal/conditional coverage (larger is faster); Values in red indicate lossless acceleration. Gray rows indicate draft models (non-reasoning).
Dataset Draft Model (DM) Mar Cov. Con Cov. DM Baseline TM Baseline $\bm{S_{\text{Mar}}}$ (×) $\bm{S_{\text{Con}}}$ (×)
QwQ-32B (RL-tuned reasoning model) as Target Model
MATH100 DeepSeek-R1-Distill-Qwen-1.5B $87.0$ $94.0$ $83.0$ $96.0$ 1.76 1.19
Qwen2.5-7B-Instruction $86.0$ 96.0 $84.0$ 96.0 7.19 5.35
DeepSeek-R1-Distill-Llama-8B 96.0 96.0 $85.0$ 96.0 1.36 1.38
Llama-3.1-8B-Instruct $87.0$ $95.0$ $75.0$ $96.0$ 2.12 2.20
AIME24 DeepSeek-R1-Distill-Qwen-1.5B 86.7 $66.7$ $60.0$ 86.7 4.03 2.60
Qwen2.5-7B-Instruction $46.7$ $33.3$ $33.3$ $86.7$ 5.71 10.10
DeepSeek Llama-3.1-8B-Instruct $80.0$ $80.0$ $80.0$ $86.7$ 2.72 2.33
Llama-3.1-8B-Instruct $33.3$ $40.0$ $13.3$ $86.7$ 4.46 2.79
AIME25 DeepSeek-R1-Distill-Qwen-1.5B $53.3$ $46.7$ $40.0$ $73.3$ 2.04 1.21
Qwen2.5-7B-Instruction $33.3$ $40.0$ $26.7$ $73.3$ 14.50 12.82
DeepSeek Llama-3.1-8B-Instruct $66.7$ $60.0$ $46.7$ $73.3$ 2.06 1.77
Llama-3.1-8B-Instruct $26.7$ $33.3$ $20.0$ $73.3$ 6.71 2.34
AMC23 DeepSeek-R1-Distill-Qwen-1.5B $88.0$ $90.0$ $74.0$ $94.0$ 1.01 1.21
Qwen2.5-7B-Instruction $76.0$ $72.0$ $68.0$ $94.0$ 10.42 8.20
DeepSeek Llama-3.1-8B-Instruct $92.0$ 94.0 $80.0$ 94.0 1.50 1.07
Llama-3.1-8B-Instruct $68.0$ $68.0$ $44.0$ $94.0$ 3.58 1.72
s1.1-32B (SFT-tuned reasoning model) as Target Model
MATH100 DeepSeek-R1-Distill-Qwen-1.5B $86.0$ $95.0$ $83.0$ $96.0$ 0.70 0.88
Qwen2.5-7B-Instruction $89.0$ 96.0 $84.0$ 96.0 2.87 2.55
DeepSeek Llama-3.1-8B-Instruct $88.0$ $95.0$ $85.0$ $96.0$ 0.79 0.82
Llama-3.1-8B-Instruct $79.0$ $85.0$ $75.0$ $96.0$ 1.33 1.56
AIME24 DeepSeek-R1-Distill-Qwen-1.5B $73.3$ $80.0$ $60.0$ $86.7$ 4.37 2.16
Qwen2.5-7B-Instruction $40.0$ $33.3$ $33.3$ $86.7$ 22.22 13.54
DeepSeek Llama-3.1-8B-Instruct $73.3$ $73.3$ $80.0$ $86.7$ 2.36 3.07
Llama-3.1-8B-Instruct $26.7$ $40.0$ $13.3$ $86.7$ 5.62 3.36
AIME25 DeepSeek-R1-Distill-Qwen-1.5B $60.0$ $60.0$ $40.0$ $66.7$ 2.87 2.03
Qwen2.5-7B-Instruction $33.3$ $40.0$ $26.7$ $66.7$ 18.52 11.90
DeepSeek Llama-3.1-8B-Instruct 66.7 $60.0$ $46.7$ 66.7 2.31 1.66
Llama-3.1-8B-Instruct $26.7$ $20.0$ $20.0$ $66.7$ 5.59 4.15
AMC23 DeepSeek-R1-Distill-Qwen-1.5B $86.0$ $86.0$ $74.0$ $96.0$ 1.37 0.98
Qwen2.5-7B-Instruction $74.0$ $78.0$ $68.0$ $96.0$ 14.49 11.36
DeepSeek Llama-3.1-8B-Instruct $92.0$ 96.0 $80.0$ 96.0 1.56 1.59
Llama-3.1-8B-Instruct $54.0$ $64.0$ $44.0$ $96.0$ 5.03 3.56
@@ 4 Experiment
We provide the hyperparameters and other task details used in our experiments in Appendix 4 and evaluate performance under two settings: marginal coverage (Mar Acc.), which measures whether the budget of prediction set align on average across test inputs, and conditional coverage (Con Acc.), which imposes a stricter requirement that the guarantee holds for each individual input instance.
@@
Table 1 reports the results of asynchronous test-time scaling when the draft model (DM) and target model (TM) come from different families. We have the following key takeaways: i) ATTS can match the performance of the target model itself. This approach effectively reduces computational overhead while maintaining high-quality outputs, up to 22.22x acceleration. ii) The most challenging datasets, AIME24/25, show strong performance in marginal coverage setup, while the other two datasets (MATH100 and AMC23) demonstrate superior results in conditional coverage setup. This highlights that while marginal coverage allocates more computational resources to the most difficult parts of the tasks effectively, conditional coverage ensures more reliable results at the individual input level, especially in simpler tasks, ensuring that each question is answered correctly. iii) It shows that while reasoning models consume more tokens during inference, using a reasoning model as the draft model provides better scaling performance than a non-reasoning model, though the non-reasoning model offers the highest acceleration. iv) When the average length of the reasoning chain output by the draft model exceeds that of the target model, the acceleration ratio is typically less than 1 on simpler datasets such as MATH and AMC23.
Table 2: The comparison when the draft and target models are from the same family. $\bm{T^{m}_{\text{Think}}}$, $\bm{T^{m}_{\text{Reason}}}$, $\bm{T^{c}_{\text{Think}}}$, and $\bm{T^{c}_{\text{Reason}}}$ denote the token-consumption ratios (×) under the marginal coverage and conditional coverage settings, relative to the SpecThink and SpecReason baselines, respectively. Token consumption is measured in tokens, where smaller values indicate higher efficiency.
@@
Token Consumption (×)
$\bm{T^{m}_{\text{Think}}}$
$\bm{T^{m}_{\text{Reason}}}$
$\bm{T^{c}_{\text{Think}}}$
$\bm{T^{c}_{\text{Reason}}}$
MATH100 Qwen2.5-7B/32B-Instruct 86.0 81.0 79.0 73.7 0.60 0.66 0.68 0.72
s1.1-7B/32B 88.0 87.0 85.0 73.7 0.50 0.54 0.57 0.44
DeepSeek-R1-Distill-Qwen-1.5B/32B 88.0 87.0 84.0 76.7 0.48 0.52 0.53 0.39
Skywork-OR1-7B/32B 88.0 89.0 70.0 75.8 0.42 0.37 0.28 0.31
AIME24 Qwen2.5-7B/32B-Instruct 33.3 40.0 33.3 33.3 0.61 0.67 0.69 0.73
s1.1-7B/32B 66.7 73.3 40.0 26.7 0.47 0.52 0.62 0.50
DeepSeek-R1-Distill-Qwen-1.5B/32B 86.7 80.0 66.7 66.7 0.46 0.50 0.56 0.42
Skywork-OR1-7B/32B 86.7 80.0 60.0 73.3 0.33 0.27 0.19 0.24
AIME25 Qwen2.5-7B/32B-Instruct 40.0 33.3 26.7 40.0 0.62 0.68 0.70 0.74
s1.1-7B/32B 53.3 53.3 33.3 40.0 0.49 0.53 0.64 0.52
DeepSeek-R1-Distill-Qwen-1.5B/32B 60.0 53.3 46.7 35.7 0.47 0.50 0.55 0.43
Skywork-OR1-7B/32B 60.0 53.3 40.0 53.3 0.41 0.36 0.29 0.22
AMC23 Qwen2.5-7B/32B-Instruct 72.0 70.0 74.0 72.0 0.63 0.69 0.71 0.75
s1.1-7B/32B 82.0 78.0 76.0 78.0 0.48 0.52 0.62 0.48
DeepSeek-R1-Distill-Qwen-1.5B/32B 92.0 88.0 82.0 80.0 0.46 0.50 0.57 0.44
Skywork-OR1-7B/32B 96.0 94.0 82.0 86.0 0.39 0.34 0.37 0.28
@@ 4.2 Performance of Budget prediction
In this section, we evaluate the accuracy of budget prediction under marginal and conditional coverage settings. This shows how well our method controls target model interventions in rejection sampling, reflecting the accuracy of conformal prediction in estimating the rejection rate.
@@
In Figure 6(b), we report the accuracy of the target-model intervention rate under marginal coverage, where the rejection rate is predicted at the dataset level. Budget-prediction accuracy across the full dataset is high, especially with the 64-sample configuration, whose absolute error stays within $5\%$. With $K_{d}=500$ tokens for calibration and $K_{d}=500$ for sampling, the error remains within $2\%$. This directly highlights the importance of constructing a diverse calibration set for maintaining high prediction accuracy.
@@
In practice, we require precise per-batch budget control, rather than a single aggregate budget over the entire test set. Due to the limited capacity of the target model’s server, it cannot process all requests concurrently. As a result, inference is performed in batches, with the token budget enforced for each batch to meet the load constraint. In Figure 6(a), we report the accuracy of the target model intervention rate under conditional coverage. Under online calibration with a rejection rate of $25\%$, when $K_{d}=500$ in both the calibration and sampling stages, the 16-sample and 64-sample settings achieve similar accuracy. However, when the calibration stage uses $K_{d}=500$ but the sampling stage uses $K_{d}=700$, the 64-sample setting attains significantly higher budget-prediction accuracy. This indicates that increasing the number of parallel samples can improve budget prediction accuracy when the sampling token budget differs from the calibration token budget (i.e., under a calibration–sampling token budget mismatch).
@@ 4.3 Asynchronous Test-time Scaling within The Same Model Families
In this section, we examine the performance across models within the same family, including both reasoning and non-reasoning models in Table 2. In this setting, since the target model and draft models share the same vocabulary, we can compare against baselines that are only applicable to models within the same family, such as Speculative Thinking (Yang et al., 2025), denoted as SpecThink. Our findings are as follows: i): When the draft and target models belong to the same family, in most cases, the best performance is achieved under the setting of marginal coverage, even on simpler datasets like MATH and AMC23. ii): Across datasets, DeepSeek and Skywork show the strongest gains on challenging benchmarks (AIME, AMC), while Qwen2.5 performs competitively on MATH100 but lags significantly on harder tasks. iii): Moreover, s1.1 achieves moderate improvements, usually surpassing Qwen but not reaching the level of Skywork or DeepSeek. iv): Finally, SpecReason and SpecThink generally underperform compared with ATTS and consume more tokens, especially when the draft model is a reasoning model or on the more challenging AIME dataset, suggesting that their effectiveness remains limited on more complex reasoning tasks.
Figure 7: Token Consumption with 16 Sampling.
@@ 4.4 Analysis of token budget and latency
In this part, we analyze the trade-offs between latency and throughput under different sampling and token budget settings. i): As shown in Figure 10(a), the latency of SpecReason and SpecThink consistently increases with the number of samples, highlighting the cost of scaling up sampling. In contrast, our method significantly reduces the sampling latency and achieves the lowest inference latency under the condition coverage setting. ii): The time overhead of online calibration is nearly negligible, particularly at larger sample sizes. iii): Meanwhile, Figure 10(b) illustrates how throughput varies with the per-sample token budget for methods such as SpecReason and SpecThink, revealing diminishing returns as the token budget becomes large. Our method is able to maintain high throughput even under very large budgets, especially in the marginal coverage setting. Overall, these results (Figure 8) provide insights into the balance between efficiency and performance when designing inference strategies. Under the setting of 16 samples and no limit on the maximum sequence scaling turns per sample, we achieved a 56.7x speedup in inference and a 4.14x throughput improvement compared to the baseline.
@@
Figure 7 presents the token consumption under the 16-sample setting. This includes sampling performed solely by the target model or the draft model as baselines, as well as asynchronous sampling under both condition coverage and marginal coverage settings. Compared with the two baselines, our method can significantly reduce token consumption, especially under the condition coverage setting, as it enables budget prediction at the instance level.
@@
We presented ATTS (Asynchronous test-time scaling), a framework that addresses the core inefficiencies of test-time scaling in LLMs. By refining arithmetic intensity and introducing online calibration with a rejection sampling pipeline, ATTS effectively controls rejection rates while reducing latency and memory overhead. Experiments on multiple reasoning benchmarks confirm that ATTS achieves better efficiency and reliability than speculative baselines. This work establishes ATTS as a practical and principled approach for scalable test-time scaling, with potential extensions to dynamic adaptation and real-world deployment.